An rms voltage of 22.8 V with a frequency of 1.50 kHz is applied to a 0.250 mH inductor.

A.What is the rms current in this circuit? (In Amperes)

B.By what factor does the current change if the frequency of the voltage is doubled?
Answer #1

Impedence = Inductive reactance XL = wL = 2pif L

XL = 2* 3.14 * 1500 * 0.25 e -3

XL = 2.355 Ohms


SO Current Irms = Vrms/XL = 22.8/2.355

Irms = 9.68 Amps

------------------------------------------------------

if f doubled wL also doubled

if V is doubled

then Irms 2V/2xL = 9.68 Amps

so remains same